OpenStudy (anonymous):

my next question is 4x^2 cube root of 3 times 2x cube root of 3

OpenStudy (anonymous):

\[(4x ^{2}\sqrt[3]{3}) (2x \sqrt[3]{3})=8x ^{3}\sqrt[3]{9}\]

OpenStudy (anonymous):

Did I interpret that correctly?